# Mixed Quantum–Classical Dynamics Yields Anharmonic Rabi Oscillations

### v1

@@ III Mixed quantum–classical representation
while an average over 100 000 trajectories is taken, which proved necessary to ensure convergence. As is obvious from Fig. 3, such Wigner sampling performs poorly in reproducing the Rabi oscillations, yielding rapidly-decaying population dynamics instead of persistent oscillations. From Fig. 2, the rapid decay is attributed to Wigner sampling effectively taking an average over oscillatory dynamics with different $n_{0}$ values for which the distribution of frequencies is broad. In this process, $P_{\mathrm{e}}$ equilibrates to a value exceeding 0.5 in Fig. 3, attributable to the population bias towards the excited state of the TLS, resulting from the anharmonic term in the Duffing equation, Eq. 30.
The preference of enforcing $n_{0}$ to (roughly) match the zero-point energy, instead of Wigner sampling of trajectories, resonates with previous studies. At times referred to as “focused sampling”, such reinforcement of physical properties at the single-trajectory level Stock and Müller (1999) is commonly envisioned as a means to accelerate convergence with respect to the number of trajectories Bonella and Coker (2003); Bonella, Montemayor, and Coker (2005); Kim, Nassimi, and Kapral (2008) and has found applications in conjunction with various quasiclassical approaches Mandal, Yamijala, and Huo (2018); Runeson and Richardson (2019); Ananth (2022). Importantly, while focused sampling enables convergence at the single-trajectory level for the Rabi model studied here, its most essential effect is the remedying of unphysical behaviors arising under Wigner sampling. It should be noted that unphysical behaviors under Wigner sampling have been observed in previous studies Tempelaar and Reichman (2017); Bondarenko and Tempelaar (2023); Hsieh, Krotz, and Tempelaar (2023b).
@@ IV Discussion
In summary, we have shown that Rabi oscillations can to a good approximation be reproduced by MQC dynamics, even in the single-quantum limit. The resulting population dynamics is governed by an undamped and unforced Duffing equation, Eq. 30, involving an anharmonic term, the solutions of which consist of persistent oscillatory functions with frequencies dependent on the initial occupancy of the optical mode. Rough agreement with the quantum result is found when this occupancy is initialized such that the zero-point energy is approximately enforced for the optical mode. These results underscore the advantage of so-called focused sampling techniques Stock and Müller (1999) in MQC dynamics over commonly-applied Wigner sampling. However, which approach is preferred may be highly susceptible to the problem at hand; something that is worthy of further inquiry.
@@
As mentioned before, the quantum Rabi model has previously been subjected to a self-consistent MQC treatment such as presented here, but within a Pauli matrix representation A. Cives-esclop and Sánchez-Soto (1999). This analysis resulted in 5 coupled differential equations, upon which solutions could be derived in the form of Jacobi elliptic functions. As such, these solutions shows consistency with our Duffing equation analysis. We note that the corresponding dynamics was previously assessed exclusively for larger quantum numbers, for which the Jacobi elliptic functions become harmonic.
@@
Here, we restricted ourselves to the scenario of a TLS interacting with a single optical mode. Previous studies have considered the performance of MQC dynamics under Wigner sampling to the scenario where the number of optical modes is large Hoffmann et al. (2019a, b); Saller, Kelly, and Geva (2021); Hsieh, Krotz, and Tempelaar (2023a, b). Under application of the Ehrenfest theorem, Eq. 22, this approach was shown to yield quantitative inaccuracies Hoffmann et al. (2019a, b); Saller, Kelly, and Geva (2021), and at times qualitative failures Hsieh, Krotz, and Tempelaar (2023a, b). These shortcomings are remedied by a recently-proposed scheme wherein optical vacuum fluctuations are selectively decoupled from the atomic ground state Hsieh, Krotz, and Tempelaar (2023a, b). It will thus be of interest to revisit such a “decoupled” Ehrenfest approach and incorporate the focused sampling scheme adopted in the present work, in order to deliver a MQC dynamics method providing optimal accuracy for both small and large mode numbers. Lastly, it is worth noting that the Rabi model bears similarities with other canonical scenarios, including the spin–boson model. As such, our findings may offer guidance to the application of MQC dynamics beyond the quantum Rabi model. Altogether, these efforts open new realms of accurate MQC modeling.
@@ Acknowledgement
The authors thank Alex Krotz, Ken Miyazaki, and Antonio Garzón-Ramírez for helpful discussions. Research reported in this publication was supported, in part, by the International Institute for Nanotechnology at Northwestern University. M.-H.H. gratefully acknowledges support from the Ryan Fellowship and the International Institute for Nanotechnology at Northwestern University.

### v2

@@ III Mixed quantum–classical representation
while an average over 100 000 trajectories is taken, which proved necessary to ensure convergence.
As is obvious from Fig. 3, such Wigner sampling performs poorly in reproducing the Rabi oscillations, yielding rapidly-decaying population dynamics instead of persistent oscillations. From Fig. 2, the rapid decay is attributed to Wigner sampling effectively taking an average over oscillatory dynamics with different $n_{0}$ values for which the distribution of frequencies is broad, leading the resulting oscillations to destructively interfere. In this process, $P_{\mathrm{e}}$ equilibrates to a value exceeding 0.5 in Fig. 3, due to the aforementioned population bias towards the excited state of the TLS induced by the anharmonic term in the Duffing equation, Eq. 30.
The preference of enforcing $n_{0}$ to (roughly) match the zero-point energy, instead of a Wigner sampling of trajectories, resonates with previous studies. At times referred to as “focused sampling”, such reinforcement of physical properties at the single-trajectory level Stock and Müller (1999) is commonly envisioned as a means to accelerate convergence with respect to the number of trajectories Bonella and Coker (2003); Bonella, Montemayor, and Coker (2005); Kim, Nassimi, and Kapral (2008) and has found applications in conjunction with various quasiclassical approaches Mandal, Yamijala, and Huo (2018); Runeson and Richardson (2019); Ananth (2022). Importantly, while focused sampling enables convergence at the single-trajectory level for the Rabi model studied here, its most essential effect is the remedying of unphysical behaviors arising under Wigner sampling. It should be noted that unphysical behaviors under Wigner sampling have been observed in previous studies Tempelaar and Reichman (2017); Bondarenko and Tempelaar (2023); Hsieh, Krotz, and Tempelaar (2023b). Lastly, we should emphasize that focused sampling and Wigner sampling differ exclusively by the initialization of the classical coordinates, leaving the equations of motion identical.
@@ IV Discussion
In summary, we have shown that Rabi oscillations can to a good approximation be reproduced by MQC dynamics, even in the single-quantum limit. Upon adopting the RWA, the resulting population dynamics is governed by an undamped and unforced Duffing equation, Eq. 30, involving an anharmonic term, the solutions of which consist of persistent oscillatory functions with frequencies dependent on the initial occupancy of the optical mode. Rough agreement with the quantum result is found for the case of a TLS initiated in its excited state when approximately enforcing zero-point energy for the initial optical occupancy. We note that similar agreement is borne out upon raising this occupancy by a quantum while initializing the TLS in its ground state (see the Supplementary Material). These results underscore the advantage presented by so-called focused sampling techniques Stock and Müller (1999) in MQC dynamics over commonly-applied Wigner sampling. However, which approach is preferred may be highly susceptible to the problem at hand; something that is worthy of further inquiry.
@@
The dynamics of the quantum Rabi model has previously been evaluated based on a self-consistent MQC approach such as presented here, but within a Pauli matrix representation A. Cives-Esclop and Sánchez-Soto (1999). This analysis resulted in 5 coupled differential equations, upon which solutions could be derived in the form of Jacobi elliptic functions. As such, these solutions show consistency with our analysis which instead reduces to a single differential equation. We note that the corresponding dynamics was previously assessed exclusively for larger quantum numbers A. Cives-Esclop and Sánchez-Soto (1999), for which the Jacobi elliptic functions become harmonic. Indeed, within the Duffing equation, we find the anharmonicity to lose significance with increasing quantum numbers, while at the same time the exact value of the optical occupancy adopted to represent zero-point energy becomes less critical. We note that the limit of large quantum numbers could be representative of laser-driven systems, and marks a stark contrast with the quantum-optical limit explored in our work.
@@
Whereas we restricted ourselves to the scenario of a TLS interacting with a single optical mode, previous studies have considered the performance of MQC dynamics under Wigner sampling for the scenario where the number of optical modes is large Hoffmann et al. (2019a, b); Saller, Kelly, and Geva (2021); Hsieh, Krotz, and Tempelaar (2023a, b). Under application of the Ehrenfest theorem, Eq. 22, MQC dynamics was shown to yield quantitative inaccuracies Hoffmann et al. (2019a, b); Saller, Kelly, and Geva (2021), and at times qualitative failures Hsieh, Krotz, and Tempelaar (2023a, b). These shortcomings are remedied by a recently-proposed scheme wherein optical vacuum fluctuations are selectively decoupled from the atomic ground state Hsieh, Krotz, and Tempelaar (2023a, b). It will thus be of interest to revisit such a “decoupled” Ehrenfest approach and incorporate the focused sampling scheme adopted in the present work, in order to deliver a MQC dynamics method providing optimal accuracy for both small and large mode numbers.
Within the multi-modal case, nondegeneracies inevitably arise between the quantum transition energies and the classical mode frequencies. In the present work, we restricted our analysis to resonance conditions, as obtaining analytical expressions beyond resonance is nontrivial. However, in the Supplementary Material we compare transient excited-state populations produced numerically through MQC dynamics with full-quantum results, and find good agreement to be retained away from resonance.
Lastly, it is worth noting that the Rabi model bears similarities with other canonical scenarios, including the spin–boson model. Among other things, it will be of particular interest to extend our assessment of the applicability of MQC dynamics to systems involving couplings with increasing degrees of anharmonicities and non-linearity, following successful examples within the realm of quasi-classical modeling Wu and Cao (2001); Kryvohuz and Cao (2005). Such efforts will guide the application of MQC dynamics to an increasing range of phenomena.
Supplementary material
Numerical evaluation of the off-resonant quantum Rabi model and results for a ground-state initialization of the TLS.
@@ Acknowledgement
The authors thank Alex Krotz, Ken Miyazaki, and Antonio Garzón-Ramírez for helpful discussions. This material is based upon work supported by the National Science Foundation under Grant No. 2145433. Research reported in this publication was supported, in part, by the International Institute for Nanotechnology at Northwestern University. M.-H.H. gratefully acknowledges support from the Ryan Fellowship and the International Institute for Nanotechnology at Northwestern University.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.